Consider the reaction below

Al2S3(s) + H2O(l) → Al(OH)3(s) + H2S(g)

If 15.0g of aluminum sulfide and 10.0g of water are allowed to react as above, and assuming a complete reaction
a. by calculation, find out which is the limiting reagent.
b. calculate the maximum mass of H2S which can be formed from these reagents.
c. calculate the mass of excess reagent remaining after the reaction is complete.
Answer #1

a)

Molar mass of Al2S3,

MM = 2*MM(Al) + 3*MM(S)

= 2*26.98 + 3*32.07

= 150.17 g/mol

mass(Al2S3)= 15.0 g

use:

number of mol of Al2S3,

n = mass of Al2S3/molar mass of Al2S3

=(15 g)/(1.502*10^2 g/mol)

= 9.989*10^-2 mol

Molar mass of H2O,

MM = 2*MM(H) + 1*MM(O)

= 2*1.008 + 1*16.0

= 18.016 g/mol

mass(H2O)= 10.0 g

use:

number of mol of H2O,

n = mass of H2O/molar mass of H2O

=(10 g)/(18.02 g/mol)

= 0.5551 mol

Balanced chemical equation is:

Al2S3 + 6 H2O ---> 3 H2S + 2 Al(OH)3

1 mol of Al2S3 reacts with 6 mol of H2O

for 9.989*10^-2 mol of Al2S3, 0.5993 mol of H2O is required

But we have 0.5551 mol of H2O

so, H2O is limiting reagent

Answer: H2O

b)

Molar mass of H2S,

MM = 2*MM(H) + 1*MM(S)

= 2*1.008 + 1*32.07

= 34.086 g/mol

According to balanced equation

mol of H2S formed = (3/6)* moles of H2O

= (3/6)*0.5551

= 0.2775 mol

use:

mass of H2S = number of mol * molar mass

= 0.2775*34.09

= 9.46 g

Answer: 9.46 g

C)

According to balanced equation

mol of Al2S3 reacted = (1/6)* moles of H2O

= (1/6)*0.5551

= 9.251*10^-2 mol

mol of Al2S3 remaining = mol initially present - mol reacted

mol of Al2S3 remaining = 9.989*10^-2 - 9.251*10^-2

mol of Al2S3 remaining = 7.376*10^-3 mol

Molar mass of Al2S3,

MM = 2*MM(Al) + 3*MM(S)

= 2*26.98 + 3*32.07

= 150.17 g/mol

use:

mass of Al2S3,

m = number of mol * molar mass

= 7.376*10^-3 mol * 1.502*10^2 g/mol

= 1.108 g

Answer: 1.11 g